solve algebra problem - the attendance at a ball fame was 400 people. students tickets cost $2 and adults tickets cost $3. if $1050 was collected in ticket sales how many if each type were sold.
Answer by macston(5194)

.
A=number of adult tickets; S=number of student tickets
.
A+S=400
S=400-A
.
$2S+$3A=$1050
$2(400-A)+$3A=$1050
$800-$2A+$3A=$1050
$1A=$250
A=250 ANSWER 1: 250 adult tickets were sold.
S=400-A=400-250=150 ANSWER 2: 150 student tickets were sold.
.
CHECK:
.
$2S+$3A=$1050
$2(150)+$3(250)=$1050
$300+$750+$1050
$1050=$1050